# Quantum Hamiltonian Descent for Graph Partition

### v2

@@ V-C Evaluation on Small Size Networks
Our experimental evaluation compared the Quantum Hamiltonian Descent approach with the GUROBI-based modularity optimization across a diverse set of network instances. As shown in the Table I the test bed has 10 networks ranging from graph with (52 nodes, 146 edges) to instance of (1,034 nodes, 26,749 edges), with edge densities from 3.4% to 15.2%. As shown in the Figure 3, the QHD-based method demonstrated robust performance, achieving higher modularity scores in 8 out of 10 instances compared to the GUROBI implementation. The average modularity difference of 0.0029 in favor of QHD suggests that quantum-inspired optimization can effectively match or slightly exceed traditional exact optimization approaches. This robust performance, where QHD achieves higher modularity scores in 8 out of 10 test instances while using only 20% of GUROBI’s computational time with four GPUs, demonstrates its practical advantages over classical modularity optimization methods, with the potential for even greater efficiency gains through increased GPU parallelization.
@@ V-D Evaluation on Large Size Networks
The Table II and Figure 4 of GUROBI and QHD performance across different networks reveal a significant density-dependent pattern. In the Facebook network (density 0.0108), QHD achieves a notable 5.49% higher modularity score (0.7512 ± 0.0258) compared to GUROBI (0.7121 ± 0.0579), while also demonstrating better stability with lower variance. The densest network, Chameleon (density 0.0121), shows nearly identical performance between the methods, with only a 0.19% difference favoring GUROBI. The TVShow network, despite its lower density (0.0023), exhibits near performance from both methods, with QHD showing a marginal advantage of 0.33% (QHD: 0.8223 ± 0.0025, GUROBI: 0.8196 ± 0.0044) and both methods maintaining very low variance. Notably, in the sparsest network, LastFM (density 0.0010), GUROBI outperforms QHD by 3.79%. These results indicate that QHD’s advantages are most pronounced in moderately dense networks. And GUROBI maintains good performance across network types, particularly in sparser networks, which is attributed to its branch-and-cut algorithm.
@@ VI Conclusion
This work demonstrates the effectiveness of quantum-inspired optimization for graph partition through our QHD-based approach. By reformulating graph partition as a QUBO problem and leveraging GPU-accelerated quantum-inspired optimization, we achieve comparable or superior performance to traditional methods while reducing computational time. Our experimental results show strong performance on moderately dense networks, with up to 5.49% improvement in modularity scores and enhanced stability compared to GUROBI. The implementation of multi-GPU acceleration suggests further scalability potential for larger networks. These findings highlight the promising intersection of quantum-inspired algorithms and high-performance computing for complex network analysis tasks. Future work could explore extending this approach to other graph optimization problems and investigating additional acceleration techniques for handling ultra-large-scale networks. Better designed algorithms to formulate graphs into denser matrices can reduce the number of variables in QUBO, and combination with high-performance sparsity computation will also be helpful. Our results demonstrate that quantum-inspired methods, combined with modern computing architectures, offer a practical pathway for advancing the capabilities of graph partition in real-world network analysis applications.

### v3

@@ V-C Evaluation on Small Size Networks
Figure 4: Solution Quality Comparison When GUROBI Reaches Optimality: Among 199 instances where GUROBI found optimal solutions (predominantly in smaller problems), QHD achieved identical solutions in 75.4% of cases. In cases where QHD did not match GUROBI’s optimal solutions (24.6% of optimally solved instances), the objective value differences remained within a 1.6% relative gap, primarily occurring in smaller instances where the absolute differences in objective values were minimal.
@@
Our experimental evaluation compared the Quantum Hamiltonian Descent approach with the GUROBI-based modularity optimization across a diverse set of network instances. As shown in Table I, the test bed has 10 networks ranging from a graph with (52 nodes and 146 edges) to an instance of (1,034 nodes, 26,749 edges), with edge densities from 3.4% to 15.2%. As shown in Fig. 5, the QHD-based method demonstrated robust performance, achieving higher modularity scores in 8 out of 10 instances compared to the GUROBI implementation. The average modularity difference of 0.0029 in favor of QHD suggests that quantum-inspired optimization can effectively match or slightly exceed traditional exact optimization approaches. This robust performance, where QHD achieves higher modularity scores in 8 out of 10 test instances while using only 20% of GUROBI’s computational time with four GPUs, demonstrates its practical advantages over classical modularity optimization methods, with the potential for even greater efficiency gains through increased GPU parallelization.
@@ V-D Evaluation on Large Size Networks
Table II and Figure 6 of GUROBI and QHD performance across different networks reveal a significant density-dependent pattern. In the Facebook network (density 0.0108), QHD achieves a notable 5.49% higher modularity score (0.7512 ± 0.0258) compared to GUROBI (0.7121 ± 0.0579), while also demonstrating better stability with lower variance. The densest network, Chameleon (density 0.0121), shows nearly identical performance between the methods, with only a 0.19% difference favoring GUROBI. The TVShow network, despite its lower density (0.0023), exhibits near performance from both methods, with QHD showing a marginal advantage of 0.33% (QHD: 0.8223 ± 0.0025, GUROBI: 0.8196 ± 0.0044) and both methods maintaining very low variance. Notably, in the sparsest network, LastFM (density 0.0010), GUROBI outperforms QHD by 3.79%. These results indicate that QHD’s advantages are most pronounced in moderately dense networks. And GUROBI maintains good performance across network types, particularly in sparser networks, which is attributed to its branch-and-cut algorithm.
@@ VI Conclusion
This work demonstrates the effectiveness of quantum-inspired optimization for community detection through our QHD-based approach. By reformulating community detection as a QUBO problem and leveraging GPU-accelerated quantum-inspired optimization, we achieve comparable or superior performance to traditional methods while reducing computational time. Our experimental results show strong performance on moderately dense networks, with up to 5.49% improvement in modularity scores and enhanced stability compared to GUROBI. The implementation of multi-GPU acceleration suggests further scalability potential for larger networks. These findings highlight the promising intersection of quantum-inspired algorithms and high-performance computing for complex network analysis tasks. Future work could explore extending this approach to other graph optimization problems and investigating additional acceleration techniques for handling ultra-large-scale networks. Better designed algorithms to formulate graphs into denser matrices can reduce the number of variables in QUBO, and combination with high-performance sparsity computation will also be helpful. Our results demonstrate that quantum-inspired methods, combined with modern computing architectures, offer a practical pathway for advancing the capabilities of community detection in real-world network analysis applications.
VII Ackowledgement
JL is supported in part by the University of Pittsburgh, School of Computing and Information, Department of Computer Science, Pitt Cyber, PQI Community Collaboration Awards. And by NASA under award number 80NSSC25M7057. This research used resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725. RL is supported by NSF Grants 2322919, 2420632, 2426031, 2426940, 2114113, and DOE Grant DE-SC0022069. We thank Yuxiang Peng from University of Maryland and Jiaqi Leng from UC Berkely for their helpful and insightul discussions.